(Yz+2x)dx+(zx+2y)dy+(xy+2z)dz=0
1
Expert's answer
2020-04-29T16:17:46-0400

We can rewrite the equation: "(yzdx+zxdy+xydz)+(2xdx+2ydy+2zdz)=0"

"yzdx+zxdy+xydz=d(xyz)"

"2xdx+2ydy+2zdz=d(x^2+y^2+z^2)"

Then "d(x^2+y^2+z^2+xyz)=0"

So "x^2+y^2+z^2+xyz=C"

Answer: "x^2+y^2+z^2+xyz=C"